How many molecules (not moles) of NH_3 are produce
Answer #1

Balanced equation:
N2 + 3 H2 = 2 NH3

8.5 x 10- 4 gm of H2 = 8.5 x 10- 4 / 2 = 0.0004216 Moles

0.0004216 of H2 can produce 0.0002811 Moles of NH3

Number of ammonia Molecule = 0.0002811 x 6.023 x 1023 = 1.693 x 1020 Molecules

1.693 x 1020 Molecules can be produced with 8.5 x 10- 4 gm of H2